Determine the amount of energy, in joules, stored in a defibrillator’s 17.0 μF capacitor that has a potential difference of 9.25 ✕ 105 V.


1
Expert's answer
2022-03-28T14:28:28-0400

Answer


the amount of energy, in joules


"U=\frac{CV^2}{2}\\=\frac{17*10^{-6}*9.25*9.25*10^{10}}{2}\\=727.28*10^4J"